OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x - 2y = 14 x + 3y = 9

OpenStudy (anonymous):

part 1 : as its impossible for 1 to be equal to 3.... so its impossible for system of equation to have solutions . :D

OpenStudy (anonymous):

part 2 ?

OpenStudy (anonymous):

wait for some moment... m eating dinner :P

OpenStudy (anonymous):

lol ok..

OpenStudy (anonymous):

soo.. here it is part 2 : x - 2y = 14 x = 2y + 14 now put this value of x in second equation.. 2y + 14 + 3y = 9 5y +14 = 9 5y = 9-14 y = -5/5 y = -1 now put y = -1 in first equation x - 2(-1) = 14 x + 2 = 14 x = 12 ENJOY !! :)